A gold alloy has a density of 14.6 g/cm^3. The cross-sectionsl area of the chain is 5.00 mm^2. What length of the chain has a mass of 9.70 g?
1
Expert's answer
2018-07-06T05:38:30-0400
d = m/V

V = m/d = 9.7g/14.6g/cm3 = 0.664 cm3 = 664 mm3

Length of the chain = 664mm3/5mm2 = 132.8 mm